# Faithfulness Measurable Masked Language Models

### v2

@@ 1 Introduction
In general, the cost of proxies has been some combination of incorrect assumptions, expensive computations, or using a proxy-model (Jain & Wallace, 2019; Bastings et al., 2022; Madsen et al., 2022a). Based on previous work, we propose the following desirable, which to the best of our knowledge, no previous method satisfies in all aspects, but we satisfy:
@@
The key idea is to use the erasure-metric but without solving the out-of-distribution (OOD) issue using retraining, which avoids all the limitations. This is achieved by including masking in the fine-tuning procedure of masked language models, such that masking is always in-distribution (Figure 1(a)). This is possible because language models are heavily over-parameterized and can thus support such additional complexity. Although our approach applies to Masked Language Models (MLMs), we suspect future work could apply this idea to any language model with sufficient capacity.
Our approach is significantly different from previous literature, which is completely model agnostic. Instead, we fine-tune a model such that measuring faithfulness is easy by design. We call such designed models: inherently faithfulness measurable models (FFMs).
@@
To validate that masking is in-distribution, we generalize previous OOD detection work from computer vision (Matan et al., 2022). This serves as a statistically grounded meta-validation of the faithfulness measure itself (Figure 1(b)). Finally, once the model is validated, the erasure-metric can be applied (Figure 1(c)).
@@
However, with an FMM, measuring faithfulness is computationally cheap. Therefore, optimizing an explanation w.r.t. faithfulness is possible, as proposed by Zhou & Shah (2023). However, they did not solve the OOD issue caused by masking, as it was “orthogonal” to their idea, but our inherently faithfulness measurable model fills that gap, making it indirectly inherently explainable.
Finally, for completeness, we compare a large variety of existing explanation methods and modify some existing explanations to be able to separate positive from negative contributions. In general, we find that the explanations that take advantage of masking (occlusion-based) are more faithful than gradient-based methods. However, the robustness provided by a faithfulness measurable model also makes some gradient-based methods more faithful.
@@ 2 Related Work
Note, we do not cover the faithfulness metrics that are specific to attention (Moradi et al., 2021; Wiegreffe & Pinter, 2019; Vashishth et al., 2019), as this paper presents a faithfulness metric for importance measures in general.
@@ 2.4 Removing important information should affect the prediction
Samek et al. (2017) introduce the erasure-metric: if information (input tokens) is truly important, then removing it should result in worse model performance compared to removing random information. However, Hooker et al. (2019) argue that removing tokens introduces an out-of-distribution (OOD) issue (d).
Hooker et al. (2019) solve the ODD issue by retraining the model. They point out a limitation: the method cannot separate an unfaithful explanation from “there exist dataset redundancies”. However, this was later solved by Madsen et al. (2022a), who conclude, like Bastings et al. (2022), that faithfulness is both model and task-dependent.
@@ 3 Inherently faithfulness measurable models (FMMs)
As an alternative to existing faithfulness methods, which all aim to work with any models, we propose creating inherently faithfulness measurable models (FMMs). These models provide the typical output (e.g., classification) for a given task and, by design, provide the means to measure the faithfulness of an explanation. Importantly, this allows measuring the faithfulness of a specific model, as there is no need for proxy models, an important property in a real deployment setting.
@@ 3.3 In-distribution validation
Erasure-based metrics are only valid when the input is in-distribution, and previous works did not validate for this (Hooker et al., 2019; Madsen et al., 2022a). Additionally, in-distribution is the statistical null-hypothesis and can never be proven. However, we can validate this using an out-of-distribution (OOD) test, which we do in this paper.
@@ 6 Limitations
Another direction useful for classification tasks, is to transform CLMs into MLMs, which has been shown to be quite straightforward (Muennighoff et al., 2024). It may also be possible to simply prompt an instruction-tuned CLM, such that it understands what masking means, for example Madsen et al. (2024) prompts with “The following content may contain redacted information marked with [REDACTED]”.
@@ 7 Conclusion
To the best of our knowledge, this is the first work that proposes creating a model designed to be faithfulness measurable. Importantly, our approach is very general, simple to apply, and stratifies critical desirables that previous measures didn’t. The masked fine-tuning method does not decrease performance on all 16 tested datasets while also adding in-distribution support for token masking. This is verified using statistical OOD tests.
@@
It is worth considering the significance of this. While our proposed model is not an inherently explainable model (Jacovi & Goldberg, 2020), it is indirectly inherently explainable because it provides a built-in way to measure faithfulness, which can then be optimized for. It does this without sacrificing the generality of the model, as it is still a RoBERTa model. As such, FMMs provide a new direction for interpretability, which bridges the gap between post-hoc (Madsen et al., 2022b) and inherent interpretability (Rudin, 2019). It does so by prioritizing faithfulness measures first and then the explanation, while previous directions have worked on explanation first and then measure faithfulness.
However, beam-search is just an approximative optimizer, and Leave-one-out does occasionally outperform Beam. Future work could look at better optimization methods to improve the faithfulness of such explanations.
@@
The Diabetes and Anemia datasets used in this work are based on the anonymized open-access medical database MIMIC-III (Johnson et al., 2016). These datasets are used because they are common in the faithfulness literature (Jain & Wallace, 2019; Madsen et al., 2022a). In particular, they contain many redundancies and are long-sequence datasets. The specific authors of this paper who have worked with this data have undergone the necessary HIPAA certification in order to access this data and comply with HIPAA regulations.

### v3

@@ 1 Introduction
In general, the cost of proxies has been some combination of incorrect assumptions, expensive computations, or using a proxy-model (Jain & Wallace, 2019; Bastings et al., 2022; Madsen et al., 2022a). Based on previous work, we propose the following desirable, which to the best of our knowledge, no previous faithfulness metric for importance measures satisfies in all aspects, but we satisfy:
@@
The key idea is to use the erasure-metric but without solving the out-of-distribution (OOD) issue using retraining, which avoids all the limitations. This is achieved by including masking in the fine-tuning procedure of masked language models, such that masking is always in-distribution (Figure 1(a)). This is possible because language models are heavily over-parameterized and can thus support such additional complexity. Although our approach applies to Masked Language Models (MLMs), we suspect future work could apply this idea to any language model with sufficient capacity. For example, Vafa et al. (2021) applied a similar idea to causal language models for top-k explanations.
@@
Our approach is significantly different from previous literature on importance measures, which have taken a completely model-agnostic perspective using for example retraining (Hooker et al., 2019; Madsen et al., 2022a; Amara et al., 2023). Instead, we fine-tune a model such that measuring faithfulness of importance measures is easy by design. We call such designed models: inherently faithfulness measurable models (FFMs).
To validate that masking is in-distribution, we generalize previous OOD detection work from computer vision (Matan et al., 2022). This serves as a statistically grounded meta-validation of the faithfulness measure itself (Figure 1(b)), something that previous works have not achieved (Vafa et al., 2021; Hase et al., 2021; Madsen et al., 2022a). Finally, once the model is validated, the erasure-metric (Samek et al., 2017) can be applied (Figure 1(c)).
@@
However, with an FMM, measuring faithfulness is computationally cheap. Therefore, optimizing an IM explanation w.r.t. faithfulness is possible, as proposed by Zhou & Shah (2023). However, they did not solve the OOD issue caused by masking, as it was “orthogonal” to their idea, but our inherently faithfulness measurable model fills that gap, making it indirectly inherently explainable.
Finally, for completeness, we compare a large variety of existing importance measure explanations and modify some existing explanations to be able to separate positive from negative contributions. In general, we find that the explanations that take advantage of masking (occlusion-based) are more faithful than gradient-based methods. However, the robustness provided by a faithfulness measurable model also makes some gradient-based methods more faithful.
@@ 2 Related Work
Note, we do not cover the faithfulness metrics that are specific to attention (Moradi et al., 2021; Wiegreffe & Pinter, 2019; Vashishth et al., 2019), as this paper presents a faithfulness metric for importance measures in general. Additionally, we do not cover other explanations, such as top-k explanations by Hase et al. (2021) and Vafa et al. (2021).
@@ 2.4 Removing important information should affect the prediction
Samek et al. (2017) introduce the erasure-metric: if information (input tokens) is truly important, then removing it should result in worse model performance compared to removing random information. However, Hooker et al. (2019) argue that removing tokens introduces an out-of-distribution (OOD) issue (d), and (Amara et al., 2023) have later shown this in detail.
Hooker et al. (2019) solve the ODD issue by retraining the model. They point out an important limitation; namely, the method cannot separate an unfaithful explanation from “there exist dataset redundancies”. However, this was later solved by Madsen et al. (2022a), who conclude, like Bastings et al. (2022), that faithfulness is both model and task-dependent.
@@ 3 Inherently faithfulness measurable models (FMMs)
As an alternative to existing faithfulness methods for importance measures, which all aim to work with any models, we propose creating inherently faithfulness measurable models (FMMs). These models provide the typical output (e.g., classification) for a given task and, by design, provide the means to measure the faithfulness of an explanation. Importantly, this allows measuring the faithfulness of a specific model, as there is no need for proxy models, an important property in a real deployment setting. This idea is similar to what Hase et al. (2021) and Vafa et al. (2021) proposed for top-k explanations.
@@ 3.2 Masked fine-tuning
The high-level idea is similar to Hase et al. (2021) which enabled masking support for a fixed number of tokens. However, to support a variable number of tokens we improve upon this work by sampling a masking-ratio.
@@ 3.3 In-distribution validation
Erasure-based metrics are only valid when the input masked according to the importance measure is in-distribution, and previous works did not validate for this (Hooker et al., 2019; Madsen et al., 2022a; Hase et al., 2021; Vafa et al., 2021). Additionally, in-distribution is the statistical null-hypothesis and can never be proven. However, we can validate this using an out-of-distribution (OOD) test.
@@ 6 Limitations
Another direction useful for classification tasks, is to transform CLMs into MLMs, which has been shown to be quite straightforward (Muennighoff et al., 2024). It may also be possible to simply prompt an instruction-tuned CLM, such that it understands what masking means, for example Madsen et al. (2024a) prompts with “The following content may contain redacted information marked with [REDACTED]”.
@@ 7 Conclusion
To the best of our knowledge, this is the first work that proposes creating a model designed to be faithfulness measurable. Arguably, previous work in top-k explanations (Hase et al., 2021; Vafa et al., 2021) and counterfactual explanation (Wu et al., 2021; Kaushik et al., 2020) have indirectly achieved something similar. However, their motivation was to provide explanations or robustness, not measuring faithfulness.
Importantly, our approach is very general, simple to apply, and stratifies critical desirables that previous faithfulness measures didn’t. The masked fine-tuning method does not decrease performance on all 16 tested datasets while also adding in-distribution support for token masking, which we are able to verify down to fundamental statistics using an out-of-distribution test.
@@
It is worth considering the significance of this. While our proposed model is not an inherently explainable model (Jacovi & Goldberg, 2020), it is indirectly inherently explainable because it provides a built-in way to measure faithfulness, which can then be optimized for. It does this without sacrificing the generality of the model, as it is still a RoBERTa model. As such, FMMs provide a new direction for interpretability, which bridges the gap between post-hoc (Madsen et al., 2022b) and inherent interpretability (Rudin, 2019). It does so by prioritizing faithfulness measures first and then the explanation, while previous directions have worked on explanation first and then measure faithfulness (Madsen et al., 2024b).
However, beam-search is just an approximative optimizer, which only archives perfect explanations at infinite beam-width, and Leave-one-out does occasionally outperform Beam for that reason. Future work could look at better optimization methods to improve the faithfulness of importance measures.
@@
The Diabetes and Anemia datasets used in this work are based on the anonymized open-access medical database MIMIC-III (Johnson et al., 2016). These datasets are used as they are common in the faithfulness literature (Jain & Wallace, 2019; Madsen et al., 2022a). In particular, they contain many redundancies and long-sequence inputs. The specific authors of this paper who have worked with this data have undergone the necessary HIPAA certification in order to access this data and comply with HIPAA regulations.